Dear Math Tutor,
The function f is defined by f(x). Find the two values of x for which f(x)=f(4)
I worked out f(4) = 9, then I tried x^2-3x+5=9 and I couldn't get anywhere
With gratitude,
Nats

Answer by rothauserc(4718)

f(x) = x^2 -3x +5
we know f(4) is
f(4) = 4^2 -3*4 + 5 = 9
x^2 -3x +5 = 9
subtract 9 from both sides of =
x^2 -3x -4 = 0
factor the quadratic
(x -4)(x +1) = 0
x - 4 = 0
x = 4
x + 1 = 0
x = -1
The two values of x for which f(x)=f(4) are 4, -1